NPA Corporation has introduced a new product and expects a change in EBIT of $490,000. Currently, the company has a 32% marginal tax rate. This project will produce $110,000 of depreciation per year. Additionally, the project will cause the following changes in year 1:

Without the project/With the project

Accounts receivable $33,000 $45,000

Inventory 21,000 36,000

Accounts payable 30,000 43,000

What is the project's free cash flow in year 1?
Answer #1

chane in working capital = change in receivable+change in inventory-change in payable

=(45000-33000)+(36000-21000)-(43000-30000)=14000

FCFF = EBIT*(1-tax rate)+Depreciation- Chng. In Work. Cap.
FCFF = 490000*(1-0.32)+110000-14000
FCFF = 429200